OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

First find the slope of the line. y2-y1/x2-x1 Slope = m = 1-2/2--5 = -1/7 Next: substitute values into the equation y=mx+b 2 = (-1/7)(-2)+b 2= 2/7 + b 12/7 = b So you've got: y = -1/7x + 12/7 Standard form is Ax + By = C 1/7x+y = 12/7 7(1/7x+y = 12/7) = x+7y=12

OpenStudy (anonymous):

can you explain the ax+by=c